OpenStudy (anonymous):

HELP! WILL GIVE MEDAL! --What are the possible numbers of positive, negative, and complex zeros of f(x) = -3x4 - 5x3 - x2 - 8x + 4? (a) Positive: 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0 (b) Positive: 1; negative: 3 or 1; complex: 2 or 0 (c) Positive: 3 or 1; negative: 1; complex: 2 or 0 (d) Positive: 4 or 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@uri @jdoe0001

OpenStudy (amistre64):

decartes has a rule of sign, can you define it?

OpenStudy (anonymous):

Is it the one where if its a negative or something?

OpenStudy (anonymous):

The change in signs right?

OpenStudy (amistre64):

yes, change in sign for positive x, for negative x, and then if we have doubles we can pull out they are complexes

OpenStudy (anonymous):

Theres one sign change. I don't know how to determine whether its positive, negative etc.

OpenStudy (amistre64):

-3x4 - 5x3 - x2 - 8x + 4 let x = 1, its a postive number, and the signs (operations) dont change -3 - 5 - 1 - 8 + 4 |------------|-- 0 1 there is only 1 sign change, meaning that we have at best 1 postivie root and 0 complex roots let x = -1, its a negative and the odd powers are going to change sign: (-1)^2 = 1 and (-1)^3 = -1 so we are just swapping odd degree signs since we are multiplying by -1 -3x4 - 5x3 - x2 - 8x + 4 -3x4 + 5x3 - x2 + 8x + 4 |-----|-----|---|------- 0 1 2 3 so 3 negative roots and 0 complex roots or 1 negative root and 2 complex roots

OpenStudy (amistre64):

since only 1 option has postive: 1 and thats all we get ... then its prolly the one we want

OpenStudy (amistre64):

Positive: 1; negative: 3 or 1; complex: 2 or 0 yep yep yep

OpenStudy (anonymous):

Got it! Thanks :)

OpenStudy (amistre64):

youre welcome